### v1

# MediRound: Multi-Round Entity-Level Reasoning Segmentation in Medical Images

Abstract
Despite the progress in medical image segmentation, most existing methods remain task-specific and lack interactivity. Although recent text-prompt-based segmentation approaches enhance user-driven and reasoning-based segmentation, they remain confined to single-round dialogues and fail to perform multi-round reasoning. In this work, we introduce Multi-Round Entity-Level Medical Reasoning Segmentation (MEMR-Seg), a new task that requires generating segmentation masks through multi-round queries with entity-level reasoning. To support this task, we construct MR-MedSeg, a large-scale dataset of 177K multi-round medical segmentation dialogues, featuring entity-based reasoning across rounds. Furthermore, we propose MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. To mitigate the inherent error propagation in the chain-like pipeline of multi-round segmentation, we introduce a lightweight yet effective Judgment & Correction Mechanism during model inference. Experimental results demonstrate that our method effectively tackles the MEMR-Seg task, surpassing conventional medical referring segmentation approaches. The project is available at https://github.com/Edisonhimself/MediRound.
Figure 1: A demo dialogue of our proposed MediRound. Our model can comprehend user queries that refer to the mask results from previous rounds (e.g., the Round 2 query refers to the Round 1 mask result), enabling cross-round entity-level reasoning in multi-round medical conversations. In contrast, conventional text-prompt-based medical segmentation methods struggle in this complex task.
1 Introduction
Medical image segmentation seeks to precisely delineate regions of interest, such as organs, lesions, and tissues, across diverse medical imaging modalities [41, 31, 28]. This task supports a wide range of clinical applications [43, 15, 45, 4] and represents a fundamental enabler for continuous advancements in medical research [27]. Most existing studies focus predominantly on task-specific segmentation [26, 9, 34], commonly referred to as “specialist models”. Despite their remarkable performance, these models generally suffer from limited adaptability and lack the interactive capabilities necessary for practical, real-world deployment.
Recent advancements in text-prompt-based medical image segmentation [50, 6, 11, 20] overcome this limitation by allowing users to actively guide the model via simple textual queries, thereby enabling more diverse and user-driven segmentation outcomes. Building upon these developments, several MLLM-based approaches [39, 46, 13] further enhance the interactivity and practicality by enabling reasoning-driven medical segmentation, allowing users to obtain the desired masks through implicit queries that do not necessarily contain specialized medical terminology. However, these models primarily focus on single-round medical segmentation and are incapable of supporting multi-round and continuous text-based interactions. Moreover, although they can effectively infer implicit medical queries within a single dialogue round, they fail to perform cross-round, entity-level reasoning in multi-round conversations. As illustrated in Figure 1, in practical medical segmentation scenarios, users tend to formulate new segmentation queries based on the mask results obtained from previous rounds. For example, in Conversation 2, the user’s query is based on the mask generated in Conversation 1, and the results of Conversation 2 in turn serve as a reference for Conversation 4. Such iterative questioning not only ensures the continuity and integrity of the entire conversation but also enables users to express their desired segmentation targets more efficiently, accurately, and conveniently. Unfortunately, due to the lack of multi-round entity-level reasoning capabilities, existing methods often fail to produce the expected responses when dealing with inputs that involve cross-round logical dependencies.
In this work, we define a new medical vision task, termed MEMR-Seg (Multi-Round Entity-Level Medical Reasoning Segmentation), which entails generating binary segmentation masks based on multi-round medical image queries, involving entity-level reasoning across dialogue rounds. Notably, the queries are not only multi-round in nature, but each round is also a derivative and extended inquiry based on the entity results from the previous round. To successfully perform this task, the model should possess two essential capabilities: (1) supporting multi-round dialogue and cross-round entity-level reasoning; and (2) generating accurate segmentation results in response to user queries.
Given that data scarcity represents a major bottleneck in accomplishing MEMR-Seg, we first introduce a Multi-Round Entity-Level Medical Reasoning Segmentation dataset, termed MR-MedSeg, which contains a large number of medical conversations centered on entity-level reasoning segmentation with multi-round queries and answers. We collect all metadata from the publicly available SA-Med2D-20M [48]. Subsequently, inspired by SegLLM [42], we construct multi-round question–answer pairs through a combination of manual annotation and GPT-5–based generation. As a result, the MR-MedSeg dataset comprises a large and diverse collection of 177K conversations. Notably, a key distinction between the proposed MR-MedSeg and traditional medical segmentation datasets [2, 37, 10, 35, 36, 39] lies in the fact that our dataset not only features multi-round dialogues but also incorporates entity-based reasoning logic across different dialogue rounds. This design enhances the practical utility of MR-MedSeg and facilitates more interactive medical image segmentation.
Furthermore, building upon existing reasoning-based segmentation works [40, 18, 47, 42], we introduce MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. MediRound integrates the segmentation feature representations from the reference round and the textual information from previous dialogues with the current query. These are jointly embedded into the input sequence of the inner-MLLM. This design enables the model not only to comprehend the current query but also to access the core information from the reference round and maintain awareness of the entire conversational history. Moreover, to address the inevitable issue of error accumulation inherent in multi-round segmentation (as illustrated in Figure 1, if the result of round 1 contains errors, these errors may propagate to round 2, and the resulting inaccuracies in round 2 can further affect round 4), we introduce a lightweight yet effective Judgment and Correction Mechanism during the model inference stage. Applied after the end-to-end model training phase, this mechanism helps MediRound mitigate error propagation in practical multi-round inference and effectively enhances the overall accuracy of multi-turn entity-level reasoning. Extensive experimental results demonstrate the superior performance of MediRound in both multi-round and single-round medical segmentation, while confirming the effectiveness of the Judgment & Correction Mechanism. Overall, we summarize our main contributions as follows:
  • We introduce the Multi-Round Entity-level Medical Reasoning Segmentation (MEMR-Seg) task, which requires reasoning over multi-round queries that refer to results from earlier rounds in medical images.
  • We construct the MR-MedSeg dataset, which comprises 177K multi-round medical segmentation dialogues that require multi-round reasoning.
  • We propose an effective baseline model, MediRound, and introduce a Judgment & Correction Mechanism to address error accumulation during multi-round reasoning segmentation. Extensive experiments demonstrate the superior performance of our model and the utility of the proposed mechanism.
Figure 2: Overview of MR-MedSeg. Our dataset comprises five types of medical reasoning dialogues, each characterized by a specific form of inter-instance relationship, encompassing nearly all multi-round interaction scenarios encountered in real-world medical applications.
2 Related Works
Medical Image Segmentation.   As one of the foundational architectures in medical image segmentation, the U-Net family [34, 51, 7, 14] has inspired numerous extensions and variants. Despite their strong performance on specific tasks and imaging modalities, these fully automated and task-specific models often lack the flexibility and user interactivity necessary for adaptability in practical, real-world medical environments. Text-prompt-based segmentation models address this by allowing clinicians to guide segmentation via language prompts [8, 24, 6, 50], and recent works further support complex implicit queries [39, 13, 46], improving flexibility and interactivity. However, they cannot handle continuous, multi-turn queries or perform cross-round reasoning over medical entities. In this work, we focus on multi-round entity-based medical reasoning segmentation, enabling entity-centered reasoning across dialogue rounds while generating accurate segmentation masks.
Reasoning Segmentation.   Reasoning segmentation in natural images is introduced by LISA [18], which interprets implicit textual instructions to generate segmentation masks by combining MLLMs’ vision-language understanding [21] with SAM’s segmentation capabilities [17]. More recently, several follow-up studies extend reasoning segmentation [16, 32, 29, 49]. Among them, SegLLM [42] represents the initial effort to address multi-round reasoning segmentation in natural images. Compared to natural image domains, multi-round reasoning segmentation in medical imaging may provide greater practical value. This is because such a segmentation paradigm can not only assist clinicians in performing rapid, continuous segmentation during clinical practice, but also help lay users learn medical knowledge more effectively. However, despite recent progress in medical reasoning segmentation [39, 13, 46], this aspect remains unaddressed. In this paper, building upon existing works [19, 25, 48, 42], we develop a method for multi-round reasoning in medical image segmentation.
3 MR-MedSeg
Dataset Overview.   The overview of our proposed MR-MedSeg dataset is shown in Figure 2. As illustrated in the figure, we design five representative types of interaction scenarios tailored for multi-round entity-level medical reasoning segmentation, each reflecting realistic use cases in clinical practice. These scenarios each involve distinct forms of inter-instance relationships, including organ–lesion dependency, anatomical hierarchy, organ/tissue attribute association, spatial relationship, and strong inferential reasoning relationship. Figure 2 further presents the overall composition and proportions of these scenarios in MR-MedSeg. Additional detailed descriptions of each scenario and their corresponding inter-instance characteristics are provided in the Appendix.
Figure 3: Semi-automatic pipeline for constructing MR-MedSeg dataset. The process includes three stages: entity selection, relationship generation, and template integration. The pipeline is primarily driven by manual annotation and complemented by GPT-5–based generation.
Data Construction Pipeline.   As illustrated in Figure 3, we design a semi-automatic pipeline for constructing the MR-MedSeg dataset. The details are as follows.
In Step 1, we manually select appropriate medical entities according to the five types of multi-round medical interaction scenarios illustrated in Figure 2. Specifically, we traverse the SA-Med2D-20M dataset [48] and identify sub-datasets that meet the criteria of each scenario. From these selected sub-datasets, we further manually select the medical entities that best match the five designed interaction settings. The high-quality images, masks, and labels from SA-Med2D-20M serve as the foundation for MR-MedSeg, greatly reducing the effort required for data pre-processing and annotation. Next, in Step 2, we construct diverse inter-entity relationships for the selected medical entities based on the entity relationships defined across five scenarios, employing a hybrid strategy that combines manual annotation with GPT-5–based generation. Notably, when generating relationships that are independent of image information, such as biological or anatomical attributes shared among medical entities which are generally consistent across images, we omit the reference to the image context. In contrast, when generating relationships that depend on image information, such as the relative spatial positions of anatomical structures that vary between images, we explicitly incorporate the corresponding image context to ensure both accuracy and contextual relevance. Subsequently, in Step 3, we employ GPT-5 [1] to transform each entity relationship into 50–80 diverse yet semantically equivalent templates. The corresponding medical entities are then manually integrated into these templates to construct the initial version of the multi-round medical segmentation data. Finally, we perform an additional automatic refinement process using GPT-5 to further enhance the fluency and accuracy of the generated dialogues. Moreover, we engage three inspectors with professional medical expertise to screen and filter the data in MR-MedSeg, thereby maximizing the validity and reliability of the dataset.
Data Statistics.   The MR-MedSeg dataset comprises over 177K multi-round medical segmentation dialogues that require cross-round reasoning, along with 118K images and 569K masks. We allocate 174,934 conversations to the training set, 1,270 to the validation set, and 1,273 to the test set. In addition, MR-MedSeg encompasses a comprehensive set of 168 medical entity categories, including various types of organs, lesions, and tissues. The dataset also covers 9 medical imaging modalities, further highlighting its diversity. More detailed quantitative analyses and data examples of the dataset are provided in the Appendix.
Figure 4: Overview of the MediRound framework. The figure illustrates the model’s workflow in processing the fourth-round conversation, referring to the second-round mask. The llava-med encoder $\mathcal{G}_{v}^{enc}$ consists of the vision encoder and visual projection layer from llava-med.
Figure 5: Illustration of how the Judgment & Correction Mechanism assists MediRound in multi-round reasoning. This mechanism evaluates and optimizes the quality of the [SEG] hidden layer features in each round, effectively preventing current-round errors from being propagated to later dialogues. Notably, the mechanism is not involved in the end-to-end training of MediRound and is only introduced during the evaluation process of the model.
4 Baseline Framework
We propose an effective baseline architecture for multi-round medical reasoning segmentation, termed MediRound. MediRound demonstrates the ability to perform multi-round reasoning over segmented medical entities by understanding information from preceding rounds to achieve fine-grained segmentation in the current round. To address the inherent drawback that errors in previous rounds can severely affect the performance of subsequent rounds in multi-round segmentation pipeline, we introduce a novel Judgment & Correction Mechanism (JCM). Details of MediRound and JCM follow.
Model Architecture.   Figure 4 illustrates the overall architecture of our proposed MediRound, which is composed of four core components: MedSAM [25] serving as the vision backbone with an image encoder $\mathcal{G}_{i}^{enc}$ and a mask decoder $\mathcal{G}_{i}^{dec}$; LLaVA-Med [19] functioning as the multimodal large language foundation model $\mathcal{G}_{i}$; the LLaVA-Med visual encoder $\mathcal{G}_{v}^{enc}$ for extracting and projecting image embeddings; and a linear projection acting as the bounding box encoder $\mathcal{G}_{b}^{enc}$.
Multi-Round Reasoning and Segmentation.   Inspired by existing MLLM-based (e.g., LLaVA [21]) segmentation methods [47, 16, 49] and the pioneer in general multi-round segmentation method SegLLM [42], we propose the MediRound tailored for multi-round medical reasoning segmentation. Specifically, we first extend the original LLM vocabulary by introducing a special token, [SEG], which denotes a request for segmentation output. As shown in Figure 4, three rounds of conversations have already been established. Given the Round 4 query $x_{txt}$ together with the original input image $x_{img}$ and history conversation $\hat{x}_{txt}$ (including the question–answer pairs from Rounds 1, 2, and 3), we first utilize the mask output from the reference round (Round 2) to crop the corresponding object $x_{crop}$ from the original image. The cropped region and its bounding box coordinates $x_{box}$ in the original image are used as reference information for the referred medical entity. Subsequently, the LLaVA-Med visual encoder $\mathcal{G}_{v}^{enc}$ and the bounding box encoder $\mathcal{G}_{b}^{enc}$ are employed to extract their respective features, which are then concatenated with $x_{txt}$, $\hat{x}_{txt}$ and $x_{img}$ to form the complete input. This input design enables the model not only to comprehend the current query but also to access key information from the reference round while maintaining awareness of the full conversational history. We then feed the complete input into the MLLM LLaVA-Med $\mathcal{G}_{i}$, which in turn outputs a text response $\hat{y}_{txt}$. It can be formulated as:
$\begin{aligned} \hat{y}_{txt}=&\;\mathcal{G}_{i}(x_{img},\mathcal{G}_{v}^{enc}(x_{crop}),\mathcal{G}_{b}^{enc}(x_{box}),x_{txt},\hat{x}_{txt}).\end{aligned}$ (1)
The output $\hat{y}_{txt}$ includes the special token [SEG] when the model intends to segment a specific medical entity within $x_{img}$. We then extract the last-layer embedding of the LLM corresponding to the [SEG] token, denoted as $h_{c}$, which encapsulates the rich semantic representation of the medical entity inferred by the MLLM. Finally, the dense visual features $f$, extracted from the input image $x_{img}$ by the vision backbone $\mathcal{G}_{i}^{enc}$, are fed into the decoder $\mathcal{G}_{i}^{dec}$ together with $h_{c}$ to generate the target segmentation mask $\hat{\mathbf{M}}$. The process can be formulated as:
$\begin{aligned} f=\mathcal{G}_{i}^{enc}&\;(x_{img}),\quad\hat{\mathbf{M}}=\mathcal{G}_{i}^{dec}(f,h_{c}).\end{aligned}$ (2)
Table 1: Comparison of multi-round entity-level medical reasoning segmentation performance between MediRound and prior related methods. “regular case” refers to relatively regular reasoning dialogues, whereas “hard case” denotes more challenging ones, which involve more complex reasoning and dialogues with a greater number of turns. $\mathcal{T}$ denotes whether the method is an independent model (a single model or with an extension). This table reports the overall scores across all rounds in the MR-MedSeg dataset.
Methods val test $\mathcal{T}$
overall regular case hard case
Dice gIoU cIoU Dice gIoU cIoU Dice gIoU cIoU
Human-Thinking + BiomedParse [50] 21.1 17.0 12.1 28.6 22.1 19.0 11.3 10.2 8.4
Human-Thinking + MedPLIB [12] 33.7 25.3 24.5 35.0 24.8 24.7 32.0 25.2 23.4
Human-Thinking + IMIS-Net [6] 30.5 25.9 32.2 33.7 30.0 42.6 26.4 21.0 25.3
Human-Thinking + MediSee [39] 43.0 33.5 35.4 45.6 34.5 38.6 45.1 37.1 39.7
GPT-4o [1] + IMIS-Net [6] 19.3 15.8 19.6 23.8 18.7 20.8 17.4 15.2 21.2
Gemini-2.5-Pro [38] + IMIS-Net [6] 26.9 22.6 28.0 33.6 29.5 35.2 21.9 17.0 19.2
Qwen3-VL [3] + IMIS-Net [6] 26.6 22.8 26.0 23.4 18.3 20.1 25.1 22.3 26.2
GPT-4o [1] + MediSee [39] 36.8 28.7 29.4 45.2 34.2 38.0 33.4 26.8 28.9
Gemini-2.5-Pro [38] + MediSee [39] 33.5 26.1 27.6 38.6 28.6 30.8 37.9 31.4 35.1
Qwen3-VL [3] + MediSee [39] 42.7 34.0 34.5 45.2 34.1 38.0 37.8 31.4 32.8
SegLLM-7B [42] 36.3 27.5 36.8 42.1 42.4 32.9 30.4 31.0 22.2
SegLLM-13B [42] 38.9 31.2 39.1 45.3 44.0 33.4 33.0 33.9 24.6
MediRound + READ [30] 53.9 45.1 54.6 60.3 51.4 57.6 47.2 36.8 49.1
MediRound 55.8 46.5 55.8 61.0 52.1 58.9 48.1 38.2 50.2
MediRound + JCM 58.4 49.0 58.9 63.3 54.5 60.2 50.3 40.5 53.4
Algorithm 1 Multi-Round Inference of MediRound with Judgment & Correction Mechanism
1:Image $x_{img}$; multi-round queries $\{x_{txt}^{(t)}\}_{t=1}^{T}$; $\mathcal{G}_{i}$: LLaVA-Med; $\mathcal{G}_{i}^{dec}$: MedSAM decoder; $h_{c}$: [SEG] feature; $\text{MLP}_{J}$: Quality Judgment Module; $\text{MLP}_{C}$: Correction Module; $\beta$: the threshold used by Quality Judgment Module to judge the quality of [SEG] feature
2:Predicted masks $\{\hat{\mathbf{M}}^{(t)}\}_{t=1}^{T}$
[…]
4:for $t=1$ to $T$ do
5:  $\tilde{x}_{txt}^{(t)}\leftarrow\text{IntegrateMaskInfo}(x_{txt}^{(t)},\hat{\mathbf{M}}^{(t-1)})$
6:  $h_{c}\leftarrow\mathcal{G}_{i}(x_{img},\tilde{x}_{txt}^{(t)})$
7:  $q\leftarrow\text{MLP}_{J}(h_{c})$
8:  if $q>\beta$ then
9:   $\hat{\mathbf{M}}^{(t)}\leftarrow\mathcal{G}_{i}^{dec}(h_{c})$
10:  else
11:   $h_{c}^{\prime}\leftarrow\text{MLP}_{C}(h_{c})$; $\hat{\mathbf{M}}^{(t)}\leftarrow\mathcal{G}_{i}^{dec}(h_{c}^{\prime})$
12:  end if
13:end for
Judgment & Correction Mechanism.   During the training of MediRound, we adopt teacher forcing [44] and directly use the ground truth masks to obtain the referred information. While this strategy effectively optimizes the model parameters, it introduces a discrepancy between the training and testing scenarios: during multi-round evaluation, the trained model can only rely on its own outputs from previous rounds as reference information. As illustrated in Figure 4, the mask information from the referred round provides essential contextual guidance for the current round. This dependency introduces a potential limitation: errors from earlier segmentation rounds may accumulate and propagate, ultimately degrading the accuracy of subsequent predictions during evaluation. This effect may become increasingly pronounced as the number of segmentation rounds increases, due to the progressive accumulation and amplification of errors. Therefore, we introduce a lightweight yet effective Judgment & Correction Mechanism that enables the model to transfer higher-quality masks from previous rounds to subsequent ones during evaluation, thereby mitigating the accumulation of segmentation errors. Specifically, as illustrated in Figure 5, during the MediRound evaluation process, we employ a Quality Judgment Module to assess the quality of the hidden features associated with the output [SEG] token in the current round. If the feature quality is deemed unsatisfactory, the features are then passed to the Correction Module for feature refinement; otherwise, they are directly decoded into masks. In this way, when subsequent rounds refer to the current round result, they receive higher-quality reference mask, which facilitate more coherent cross-round reasoning. To more clearly illustrate how JCM assists MediRound in performing multi-round reasoning, we present the pseudocode of the process in Algorithm 1. Notably, both the Quality Judgment and Correction Modules are implemented as lightweight MLPs. To train them, we freeze all parameters of the end-to-end trained MediRound and leverage the frozen MediRound weights to independently optimize the two modules in a fast and efficient manner. Further training details of JCM are presented in the Appendix.
Table 2: Multi-round entity-level medical reasoning segmentation performance of MediRound and prior related methods, reported for each round. The table presents the cIoU scores from rounds #2–#8 across all MR-MedSeg validation samples.
Methods Conversation Rounds
Round #2 Round #3 Round #4 Round #5 Round #6 Round #7 Round #8
Human-Thinking + BiomedParse [50] 13.3 17.9 13.9 15.9 11.5 9.4 7.8
Human-Thinking + MedPLIB [12] 19.7 21.0 21.3 19.7 20.5 25.0 34.4
Human-Thinking + IMIS-Net [6] 34.0 27.1 30.1 33.4 37.5 37.7 32.8
Human-Thinking + MediSee [39] 43.6 37.0 34.4 31.9 29.7 37.0 25.1
GPT-4o [1] + IMIS-Net [6] 25.1 20.4 18.1 17.0 11.6 12.5 18.9
Gemini-2.5-Pro [38] + IMIS-Net [6] 27.2 28.6 29.4 26.2 22.6 20.1 30.5
Qwen3-VL [3] + IMIS-Net [6] 25.6 24.1 20.3 21.8 15.2 14.8 27.6
GPT-4o [1] + MediSee [39] 37.7 33.4 27.3 25.5 14.9 21.5 20.9
Gemini-2.5-Pro [38] + MediSee [39] 30.8 30.4 29.5 26.7 31.5 21.5 33.3
Qwen3-VL [3] + MediSee [39] 37.7 33.9 30.1 25.4 20.2 34.7 29.9
SegLLM-7B [42] 35.2 39.5 27.2 30.1 34.7 37.4 34.1
SegLLM-13B [42] 36.4 42.1 32.5 33.2 35.3 39.6 36.0
MediRound + READ [30] 49.1 58.3 55.0 56.2 62.4 63.3 44.5
MediRound 50.2 60.8 55.9 58.8 63.7 64.7 46.1
MediRound + JCM 52.6 63.6 59.7 64.6 69.5 66.3 54.8
Table 3: Vanilla medical referring segmentation results.
Methods Metrics
Dice gIoU cIoU
Medical Interaction Seg. (bbox input)
MedSAM [25] 77.7 67.1 81.0
IMIS-Net [6] 77.0 67.5 87.2
Medical Referring Seg.
Grounding DINO [22] + SAM-Med2D [5] 18.4 12.5 10.9
Grounding DINO [22] + MedSAM [25] 18.5 11.0 13.5
BiomedParse [50] 24.9 20.1 14.5
MedPLIB [12] 39.1 40.6 45.2
IMIS-Net [6] 41.3 35.5 65.8
MediSee [39] 61.2 50.4 70.8
MediRound + READ [30] 60.7 48.0 65.2
MediRound 62.1 48.3 66.3
5 Experiments
5.1 Experimental Setting
Network Architecture.   For MediRound, as described in Sec. 4, we employ LLaVA-Med-v1.5-Mistral-7B [19] as the MLLM, and MedSAM [25] as the vision backbone. The $\mathcal{G}_{b}^{enc}$ is implemented as a linear projection from 4 dimensions to 4096. For JCM, the projections in the Quality Judgment Module and Correction Module are implemented as three-layer MLPs with dimensions [256, 512, 512, 1] and [256, 512, 512, 256], respectively.
Implementation Details.   MediRound is trained on the MR-MedSeg dataset using four NVIDIA RTX A6000 GPUs (48 GB memory each) for approximately 1.5 days. The batch size per device is set to 15. To enhance training efficiency, we adopt the DeepSpeed [33] engine. Optimization is performed using the AdamW [23] optimizer with a learning rate of 0.0003. Learning rate scheduling is managed by the WarmupDecayLR scheduler, with a warm-up phase spanning 100 iterations. The LoRA rank is uniformly set to 8. For training JCM, we keep the same hyperparameter settings, and it can be completed in just half a day. All training is conducted on the MR-MedSeg training set, employing a teacher forcing strategy [44].
5.2 Multi-Round Experimental Results
The results of multi-round medical reasoning segmentation are summarized in Tables 1 and 2. Table 1 reports the overall performance across all rounds, while Table 2 provides detailed scores for each individual round of interaction. Since existing text-prompt-based medical segmentation methods lack the capability for multi-round entity-level reasoning segmentation, we first incorporate human guidance to enable a broader and more equitable comparison. Specifically, humans with medical expertise assist traditional segmentation models through multi-round reasoning, allowing these models to perform the task. As shown in Tables 1 and 2, human assistance can substantially improve the model’s ability to perform multi-round segmentation. However, the overall performance remains suboptimal. This is primarily because, even with a certain level of medical knowledge, humans may still make errors when dealing with complex multi-round medical reasoning tasks, and their coordination with the model tends to be relatively rigid. In contrast, the proposed MediRound can effectively comprehend the current query while simultaneously integrating historical information and mask results from referred rounds. By achieving this capability, it demonstrates an average improvement of approximately 15% across various evaluation metrics compared to other methods. Notably, despite its lightweight design, the proposed JCM remarkably improves the model’s overall performance, particularly as the number of interaction rounds increases (refer to Table 2). This improvement is primarily attributed to the fact that JCM refines suboptimal mask outputs from earlier rounds before they are referred in subsequent ones, thereby effectively mitigating the error accumulation inherent in the chain-like pipeline of multi-round segmentation.
We further construct a hybrid paradigm for comparison, which integrates an MLLM with a medical segmentation model. Specifically, we leverage the strong text–image reasoning capabilities of MLLMs [3, 1, 38] to interpret complex multi-round queries, and subsequently feed their generated reasoning outputs into a medical segmentation model to produce visual predictions. As shown in Tables 1 and 2, despite being supported by powerful MLLMs, these combined approaches still underperform compared with MediRound. We attribute this advantage to our model’s end-to-end training on large-scale medical image–text pairs, which enables more coherent dynamic alignment between multimodal features than the two-stage combined paradigms. In addition, we include SegLLM [42]—the exploratory work on multi-round reasoning segmentation in natural images—in our comparison. The results indicate that while SegLLM performs well in natural images, it struggles when generalized to medical scenarios.
Figure 6: Qualitative comparison with different methods on medical multi-round interactive segmentation.
5.3 Vanilla Medical Referring Segmentation
To demonstrate MediRound’s effectiveness in conventional single-round referring medical segmentation, we evaluate it against several prior related methods on the SA-Med2D-20M [48] benchmark. As summarized in Table 3, our model achieves strong performance, remaining highly competitive in traditional medical image segmentation. For context, we also report the segmentation results of IMIS-Net [6] and MedSAM [25] using conventional bounding box guidance.
Table 4: Ablation study on the choices of vision and MLLM backbones. Experiments are conducted on the MR-MedSeg val set.
Settings Dice gIoU cIoU
Seg. Head Backbone
SAM 44.9 27.6 31.9
SAM-Med2D 53.9 45.3 55.5
MedSAM (LoRA) 54.6 44.9 54.4
MedSAM 55.8 46.5 55.8
MLLM Backbone
LLaVA-v1.5 (LoRA) 53.4 42.1 53.3
LLaVA-Med (Frozen) 48.0 39.2 49.6
LLaVA-Med (LoRA) 55.8 46.5 55.8
Figure 7: Ablation study on the threshold $\beta$ value in JCM.
5.4 Qualitative Analysis
We present qualitative comparisons in Figure 6. Existing methods largely fail in multi-round reasoning for medical segmentation, whereas MediRound can understand the current query and dialogue history for accurate segmentation.
5.5 Ablation Study
Choices of $\bm{\beta}$ in JCM.   We investigate the impact of the threshold $\beta$ in the Quality Judgement Module of JCM. As shown in Figure 7, the collaboration between JCM and MediRound achieves the best performance when $\bm{\beta=0.6}$. This is likely because a smaller $\beta$ causes many suboptimal results to lose the opportunity for refinement, while an excessively large $\beta$ may instead disturb some high-quality predictions. For the definition and role of $\beta$, see Algorithm 1.
Choices of the Main Components.   Table 4 summarizes the effects of different foundational MLLMs and segmentation head backbones. The results show that LLaVA-Med achieves superior performance owing to its strong multimodal alignment capability in medical contexts, while the MedSAM performs best as the segmentation head, benefiting from large-scale medical image pretraining.
6 Conclusion
This work introduces MEMR-Seg, a new task that advances medical image segmentation toward interactive, multi-round reasoning. By constructing the large-scale MR-MedSeg dataset and developing the MediRound baseline with a helpful Judgment & Correction Mechanism, this work demonstrates the feasibility and effectiveness of medical multi-round entity-level reasoning, providing insights for future research on interactive medical segmentation.