### v3

# MediRound: Multi-Round Entity-Level Reasoning Segmentation in Medical Images

Abstract
Despite recent progress in text-prompt-based medical image segmentation, these methods are limited to single-round dialogues and fail to support multi-round reasoning, which is important for medical education scenarios. In this work, we introduce Multi-Round Entity-Level Medical Reasoning Segmentation (MEMR-Seg), a new task that requires generating segmentation masks through multi-round queries with entity-level reasoning, helping learners progressively develop their understanding of medical knowledge. To support this task, we construct MR-MedSeg, a large-scale dataset of 177K multi-round medical segmentation dialogues, featuring entity-based reasoning across rounds. Furthermore, we propose MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. To mitigate the inherent error propagation within the chain-like pipeline of multi-round segmentation, we introduce a lightweight yet effective Judgment & Correction Mechanism during model inference. Experimental results demonstrate that our method effectively addresses the MEMR-Seg task and outperforms conventional medical referring segmentation methods. The project is available at https://github.com/Edisonhimself/MediRound.
Figure 1: A medical education dialogue illustrating the proposed MediRound. Our model can comprehend user queries that refer to the mask results from previous rounds (e.g., the Round 2 query refers to the Round 1 mask result), enabling cross-round entity-level reasoning in multi-round medical conversations. This iterative querying paradigm allows learners to progressively develop their understanding of medical entities. In contrast, conventional text-prompt-based medical segmentation methods struggle with such a complex and context-dependent task.
1 Introduction
Medical image segmentation seeks to precisely delineate regions of interest (e.g., organs and lesions) across diverse medical imaging modalities [41, 31, 28]. This capability underpins a broad range of downstream medical applications, including clinical analysis, medical education, and biomedical research  [43, 15, 45, 4, 27]. Most existing studies focus predominantly on task-specific segmentation [26, 9, 34], commonly referred to as “specialist models”. Despite their remarkable performance, these models generally suffer from limited adaptability and lack the interactive capabilities necessary for practical, real-world usage.
Recent advancements in text-prompt-based medical image segmentation [50, 6, 11, 20] overcome this limitation by allowing users to actively guide the model via simple textual queries, thereby enabling more diverse and user-driven segmentation outcomes. Building upon these developments, several MLLM-based approaches [39, 46, 13] further enhance the interactivity and practicality by enabling reasoning-driven medical segmentation, allowing users to obtain the desired masks through implicit queries.
However, these models primarily focus on single-round medical segmentation and are incapable of supporting multi-round and continuous text-based interactions. Although the one-shot querying paradigm can be efficient in clinical use, it presumes high-quality prompts and substantial user expertise. As a result, it is well suited to medical experts but less conducive to the learning process of non-expert users. In medical education settings, visual understanding is typically developed progressively. As illustrated in Figure 1, learners tend to formulate new segmentation queries based on the mask results obtained from previous rounds. For example, in Round 2 Conversation, the user’s query is based on the mask generated in Round 1 Conversation, and the results of Round 2 Conversation in turn serve as a reference for Round 4 Conversation. Such iterative questioning helps learners progressively develop a better visual understanding of different medical entities and grasp the relationships among them more effectively. Moreover, these follow-up queries are often relation-driven and logically constrained, reflecting how trainees learn anatomy and pathology through relational concepts rather than precise standalone descriptions, which aligns with how learning occurs in real-world medical education. Unfortunately, due to the lack of multi-round entity-level reasoning capabilities, existing methods often fail to produce the expected responses when dealing with inputs that involve cross-round logical dependencies.
In this work, we define a new medical vision task, termed MEMR-Seg (Multi-Round Entity-Level Medical Reasoning Segmentation), which entails generating binary segmentation masks based on multi-round medical image queries, involving entity-level reasoning across dialogue rounds. Notably, the queries are not only multi-round in nature, but each round is also a derivative and extended inquiry based on the entity results from the previous round. To successfully perform this task, the model should possess two essential capabilities: (1) supporting multi-round dialogue and cross-round entity-level reasoning; and (2) generating accurate segmentation results in response to user queries.
Given that data scarcity represents a major bottleneck in accomplishing MEMR-Seg, we first introduce a Multi-Round Entity-Level Medical Reasoning Segmentation dataset, termed MR-MedSeg, which contains a large number of medical conversations centered on entity-level reasoning segmentation with multi-round queries and answers. We collect all metadata from the publicly available SA-Med2D-20M [48]. Subsequently, inspired by SegLLM [42], we construct multi-round question–answer pairs through a combination of manual annotation and GPT-5-based generation. As a result, the MR-MedSeg dataset comprises a large and diverse collection of 177K conversations. Notably, a key distinction between the proposed MR-MedSeg and traditional medical segmentation datasets [2, 37, 10, 35, 36, 39] lies in the fact that our dataset not only features multi-round dialogues but also incorporates entity-based reasoning logic across different dialogue rounds. This design enhances the educational utility of MR-MedSeg and facilitates more interactive medical image segmentation.
Furthermore, building upon existing reasoning-based segmentation works [40, 18, 47, 42], we introduce MediRound, an effective baseline model designed for multi-round medical reasoning segmentation. MediRound integrates the segmentation feature representations from the reference round and the textual information from previous dialogues with the current query. These are jointly embedded into the input sequence of the inner-MLLM. This design enables the model not only to comprehend the current query but also to access the core information from the reference round while maintaining awareness of the entire conversational history. Moreover, to address the inevitable issue of error accumulation inherent in multi-round segmentation (as illustrated in Figure 1, if the result of Round 1 contains errors, these errors may propagate to Round 2, and the resulting inaccuracies in Round 2 can further affect Round 4), we introduce a lightweight yet effective Judgment & Correction Mechanism during the model inference stage. Applied after the end-to-end model training phase, this mechanism helps MediRound mitigate error propagation in practical multi-round inference and effectively enhances the overall accuracy of multi-turn entity-level reasoning. Extensive experimental results demonstrate the superior performance of MediRound in both multi-round and single-round medical segmentation, while confirming the effectiveness of the Judgment & Correction Mechanism. Overall, we summarize our main contributions as follows:
  • We introduce the Multi-Round Entity-Level Medical Reasoning Segmentation (MEMR-Seg) task, which requires reasoning over multi-round queries that refer to results from earlier rounds in medical images.
  • We construct the MR-MedSeg dataset, which comprises 177K multi-round medical segmentation dialogues that require multi-round reasoning.
  • We propose an effective baseline model, MediRound, and introduce a Judgment & Correction Mechanism to address the inherent error accumulation during multi-round reasoning segmentation. Extensive experiments demonstrate the superior performance of our model and the utility of the proposed mechanism.
Figure 2: Overview of MR-MedSeg. Our dataset comprises five types of medical reasoning dialogues, each characterized by a specific form of inter-instance relationship, encompassing nearly all multi-round interaction scenarios encountered in real-world medical applications.
2 Related Works
Medical Image Segmentation.   As one of the foundational architectures in medical image segmentation, the U-Net family [34, 51, 7, 14] has inspired numerous extensions and variants. Despite their strong performance on specific tasks and imaging modalities, these fully automated and task-specific models often lack the flexibility and user interactivity necessary for adaptability in practical, real-world medical environments. Text-prompt-based segmentation models address this by allowing clinicians to guide segmentation via language prompts [8, 24, 6, 50], and recent works further support complex implicit queries [39, 13, 46], improving flexibility and interactivity. However, they cannot handle continuous, multi-turn queries or perform cross-round reasoning over medical entities. In this work, we focus on multi-round entity-based medical reasoning segmentation, enabling entity-centered reasoning across dialogue rounds while generating accurate segmentation masks.
Reasoning Segmentation.   Reasoning segmentation in natural images is introduced by LISA [18], which interprets implicit textual instructions to generate segmentation masks by combining MLLMs’ vision-language understanding [21] with SAM’s segmentation capabilities [17]. More recently, several follow-up studies extend reasoning segmentation [16, 32, 29, 49]. Among them, SegLLM [42] represents the initial effort to address multi-round reasoning segmentation in natural images. Compared to natural image domains, multi-round reasoning segmentation in medical imaging may provide greater practical value. This is because such a segmentation paradigm can not only assist clinicians in performing rapid, continuous segmentation during clinical practice, but also help lay users learn medical knowledge more effectively. However, despite recent progress in medical reasoning segmentation [39, 13, 46], this aspect remains unaddressed. In this paper, building upon existing works [19, 25, 48, 42], we develop a method for multi-round reasoning in medical image segmentation.
3 MR-MedSeg
Dataset Overview.   The overview of our proposed MR-MedSeg dataset is shown in Figure 2. As illustrated in the figure, we design five representative types of interaction scenarios tailored for multi-round entity-level medical reasoning segmentation, each reflecting realistic use cases in medical education practice. These scenarios each involve distinct forms of inter-instance relationships, including organ–lesion dependency, anatomical hierarchy, organ/tissue attribute association, spatial relationship, and strong inferential relationship. Figure 2 further presents the overall composition and proportions of these scenarios in MR-MedSeg. Additional detailed descriptions of each scenario and their corresponding inter-instance characteristics are provided in the Appendix.
Figure 3: Semi-automatic pipeline for constructing MR-MedSeg dataset. The process includes three stages: entity selection, relationship generation, and template integration. The pipeline is primarily driven by manual annotation and complemented by GPT-5–based generation.
Data Construction Pipeline.   As illustrated in Figure 3, we design a semi-automatic pipeline for constructing the MR-MedSeg dataset. The details are as follows.
In Step 1, we manually select appropriate medical entities according to the five types of multi-round medical interaction scenarios illustrated in Figure 2. Specifically, we traverse the SA-Med2D-20M dataset [48] and identify sub-datasets that meet the criteria of each scenario. From these selected sub-datasets, we further manually select the medical entities that best match the five designed interaction settings. The high-quality images, masks, and labels from SA-Med2D-20M serve as the foundation for MR-MedSeg, greatly reducing the effort required for data pre-processing and annotation. Next, in Step 2, we construct diverse inter-entity relationships for the selected medical entities based on the entity relationships defined across five scenarios, employing a hybrid strategy that combines manual annotation with GPT-5–based generation. Notably, when generating relationships that are independent of image information, such as biological or anatomical attributes shared among medical entities which are generally consistent across images, we omit the reference to the image context. In contrast, when generating relationships that depend on image information, such as the relative spatial positions of anatomical structures that vary between images, we explicitly incorporate the corresponding image context to ensure both accuracy and contextual relevance. Subsequently, in Step 3, we employ GPT-5 [1] to transform each entity relationship into 50–80 diverse yet semantically equivalent templates. The corresponding medical entities are then manually integrated into these templates to construct the initial version of the multi-round medical segmentation data. Finally, we perform an additional automatic refinement process using GPT-5 to further enhance the fluency and accuracy of the generated dialogues. Moreover, we engage three inspectors with professional medical expertise to screen and filter the data in MR-MedSeg, thereby maximizing the validity and reliability of the dataset.
Data Statistics.   The MR-MedSeg dataset comprises over 177K multi-round medical segmentation dialogues that require cross-round reasoning, along with 118K images and 569K masks. We allocate 174,934 conversations to the training set, 1,270 to the validation set, and 1,273 to the test set. In addition, MR-MedSeg encompasses a comprehensive set of 168 medical entity categories, including various types of organs, lesions, and tissues. The dataset also covers 9 medical imaging modalities, further highlighting its diversity. More detailed quantitative analyses and data examples of the dataset are provided in the Appendix.
Figure 4: Overview of the MediRound framework. The figure illustrates the model’s workflow in processing the fourth-round conversation, referring to the second-round mask. The llava-med encoder $\mathcal{G}_{v}^{enc}$ consists of the vision encoder and visual projection layer from llava-med.
Figure 5: Illustration of how the Judgment & Correction Mechanism assists MediRound in multi-round reasoning. This mechanism evaluates and optimizes the quality of the [SEG] hidden layer features in each round, effectively preventing current-round errors from being propagated to later dialogues. Notably, the mechanism is not involved in the end-to-end training of MediRound and is only introduced during the evaluation process of the model.
4 Baseline Framework
We propose an effective baseline architecture for multi-round medical reasoning segmentation, termed MediRound. MediRound demonstrates the ability to perform multi-round reasoning over segmented medical entities by understanding information from preceding rounds to achieve fine-grained segmentation in the current round. To address the inherent drawback that errors in previous rounds can severely affect the performance of subsequent rounds in multi-round segmentation pipeline, we introduce a novel Judgment & Correction Mechanism (JCM). Details of MediRound and JCM follow.
Model Architecture.   Figure 4 illustrates the overall architecture of our proposed MediRound, which is composed of four core components: MedSAM [25] serving as the vision backbone with an image encoder $\mathcal{G}_{i}^{enc}$ and a mask decoder $\mathcal{G}_{i}^{dec}$; LLaVA-Med [19] functioning as the multimodal large language foundation model $\mathcal{G}_{i}$; the LLaVA-Med visual encoder $\mathcal{G}_{v}^{enc}$ for extracting and projecting image embeddings; and a linear projection acting as the bounding box encoder $\mathcal{G}_{b}^{enc}$.
Multi-Round Reasoning and Segmentation.   Inspired by existing MLLM-based (e.g., LLaVA [21]) segmentation methods [47, 16, 49] and the pioneer in general multi-round segmentation method SegLLM [42], we propose the MediRound tailored for multi-round medical reasoning segmentation. Specifically, we first extend the original LLM vocabulary by introducing a special token, [SEG], which denotes a request for segmentation output. As shown in Figure 4, three rounds of conversations have already been established. Given the Round 4 query $x_{txt}$ together with the original input image $x_{img}$ and history conversation $\hat{x}_{txt}$ (including the question–answer pairs from Rounds 1, 2, and 3), we first utilize the mask output from the reference round (Round 2) to crop the corresponding object $x_{crop}$ from the original image. The cropped region and its bounding box coordinates $x_{box}$ in the original image are used as reference information for the referred medical entity. Subsequently, the LLaVA-Med visual encoder $\mathcal{G}_{v}^{enc}$ and the bounding box encoder $\mathcal{G}_{b}^{enc}$ are employed to extract their respective features, which are then concatenated with $x_{txt}$, $\hat{x}_{txt}$ and $x_{img}$ to form the complete input. This input design enables the model not only to comprehend the current query but also to access key information from the reference round while maintaining awareness of the full conversational history. We then feed the complete input into the MLLM LLaVA-Med $\mathcal{G}_{i}$, which in turn outputs a text response $\hat{y}_{txt}$. It can be formulated as:
$\begin{aligned} \hat{y}_{txt}=&\;\mathcal{G}_{i}(x_{img},\mathcal{G}_{v}^{enc}(x_{crop}),\mathcal{G}_{b}^{enc}(x_{box}),x_{txt},\hat{x}_{txt}).\end{aligned}$ (1)
The output $\hat{y}_{txt}$ includes the special token [SEG] when the model intends to segment a specific medical entity within $x_{img}$. We then extract the last-layer embedding of the LLM corresponding to the [SEG] token, denoted as $h_{c}$, which encapsulates the rich semantic representation of the medical entity inferred by the MLLM. Finally, the dense visual features $f$, extracted from the input image $x_{img}$ by the vision backbone $\mathcal{G}_{i}^{enc}$, are fed into the decoder $\mathcal{G}_{i}^{dec}$ together with $h_{c}$ to generate the target segmentation mask $\hat{\mathbf{M}}$. The process can be formulated as:
$\begin{aligned} f=\mathcal{G}_{i}^{enc}&\;(x_{img}),\quad\hat{\mathbf{M}}=\mathcal{G}_{i}^{dec}(f,h_{c}).\end{aligned}$ (2)
Table 1: Comparison of multi-round entity-level medical reasoning segmentation performance between MediRound and prior related methods. “regular case” refers to relatively regular reasoning dialogues, whereas “hard case” denotes more challenging ones, which involve more complex reasoning and dialogues with a greater number of turns. $\mathcal{T}$ denotes whether the method is an independent model (a single model or with an extension). This table reports the overall scores across all rounds in the MR-MedSeg dataset.
Methods val test $\mathcal{T}$
overall regular case hard case
Dice gIoU cIoU Dice gIoU cIoU Dice gIoU cIoU
Human-Thinking + BiomedParse [50] 21.1 17.0 12.1 28.6 22.1 19.0 11.3 10.2 8.4
Human-Thinking + MedPLIB [12] 33.7 25.3 24.5 35.0 24.8 24.7 32.0 25.2 23.4
Human-Thinking + IMIS-Net [6] 30.5 25.9 32.2 33.7 30.0 42.6 26.4 21.0 25.3
Human-Thinking + MediSee [39] 43.0 33.5 35.4 45.6 34.5 38.6 45.1 37.1 39.7
GPT-4o [1] + IMIS-Net [6] 19.3 15.8 19.6 23.8 18.7 20.8 17.4 15.2 21.2
Gemini-2.5-Pro [38] + IMIS-Net [6] 26.9 22.6 28.0 33.6 29.5 35.2 21.9 17.0 19.2
Qwen3-VL [3] + IMIS-Net [6] 26.6 22.8 26.0 25.1 22.3 26.2 23.4 18.3 20.1
GPT-4o [1] + MediSee [39] 36.8 28.7 29.4 45.2 34.2 38.0 33.4 26.8 28.9
Gemini-2.5-Pro [38] + MediSee [39] 33.5 26.1 27.6 37.9 31.4 35.1 38.6 28.6 30.8
Qwen3-VL [3] + MediSee [39] 42.7 34.0 34.5 45.2 34.1 38.0 37.8 31.4 32.8
SegLLM-7B [42] 36.3 27.5 36.8 42.1 32.9 42.4 30.4 22.2 31.0
SegLLM-13B [42] 38.9 31.2 39.1 45.3 33.4 44.0 33.0 24.6 33.9
MediRound + READ [30] 53.9 45.1 54.6 60.3 51.4 57.6 47.2 36.8 49.1
MediRound 55.8 46.5 55.8 61.0 52.1 58.9 48.1 38.2 50.2
MediRound + JCM 58.4 49.0 58.9 63.3 54.5 60.2 50.3 40.5 53.4
Algorithm 1 Multi-Round Inference of MediRound with Judgment & Correction Mechanism
1:Image $x_{img}$; multi-round queries $\{x_{txt}^{(t)}\}_{t=1}^{T}$; $\mathcal{G}_{i}$: LLaVA-Med; $\mathcal{G}_{i}^{dec}$: MedSAM decoder; $h_{c}$: [SEG] feature; $\text{MLP}_{J}$: Quality Judgment Module; $\text{MLP}_{C}$: Correction Module; $\beta$: the threshold used by Quality Judgment Module to judge the quality of [SEG] feature
2:Predicted masks $\{\hat{\mathbf{M}}^{(t)}\}_{t=1}^{T}$
[…]
4:for $t=1$ to $T$ do
5:  $\tilde{x}_{txt}^{(t)}\leftarrow\text{IntegrateMaskInfo}(x_{txt}^{(t)},\hat{\mathbf{M}}^{(t-1)})$
6:  $h_{c}\leftarrow\mathcal{G}_{i}(x_{img},\tilde{x}_{txt}^{(t)})$
7:  $q\leftarrow\text{MLP}_{J}(h_{c})$
8:  if $q>\beta$ then
9:   $\hat{\mathbf{M}}^{(t)}\leftarrow\mathcal{G}_{i}^{dec}(h_{c})$
10:  else
11:   $h_{c}^{\prime}\leftarrow\text{MLP}_{C}(h_{c})$; $\hat{\mathbf{M}}^{(t)}\leftarrow\mathcal{G}_{i}^{dec}(h_{c}^{\prime})$
12:  end if
13:end for
Judgment & Correction Mechanism.   During the training of MediRound, we adopt teacher forcing [44] and directly use the ground truth masks to obtain the referred information. While this strategy effectively optimizes the model parameters, it introduces a discrepancy between the training and testing scenarios: during multi-round evaluation, the trained model can only rely on its own outputs from previous rounds as reference information. As illustrated in Figure 4, the mask information from the referred round provides essential contextual guidance for the current round. This dependency introduces a potential limitation: errors from earlier segmentation rounds may accumulate and propagate, ultimately degrading the accuracy of subsequent predictions during evaluation. This effect may become increasingly pronounced as the number of segmentation rounds increases, due to the progressive accumulation and amplification of errors. Therefore, we introduce a lightweight yet effective Judgment & Correction Mechanism that enables the model to transfer higher-quality masks from previous rounds to subsequent ones during evaluation, thereby mitigating the accumulation of segmentation errors. Specifically, as illustrated in Figure 5, during the MediRound evaluation process, we employ a Quality Judgment Module to assess the quality of the hidden features associated with the output [SEG] token in the current round. If the feature quality is deemed unsatisfactory, the features are then passed to the Correction Module for feature refinement; otherwise, they are directly decoded into masks. In this way, when subsequent rounds refer to the current round result, they receive higher-quality reference mask, which facilitate more coherent cross-round reasoning. To more clearly illustrate how JCM assists MediRound in performing multi-round reasoning, we present the pseudocode of the process in Algorithm 1. Notably, both the Quality Judgment and Correction Modules are implemented as lightweight MLPs. To train them, we freeze all parameters of the end-to-end trained MediRound and leverage the frozen MediRound weights to independently optimize the two modules in a fast and efficient manner. Further training details of JCM are presented in the Appendix.
Table 2: Multi-round entity-level medical reasoning segmentation performance of MediRound and prior related methods, reported for each round. The table presents the cIoU scores from rounds #2–#8 across all MR-MedSeg validation samples.
Methods Conversation Rounds
Round #2 Round #3 Round #4 Round #5 Round #6 Round #7 Round #8
Human-Thinking + BiomedParse [50] 13.3 17.9 13.9 15.9 11.5 9.4 7.8
Human-Thinking + MedPLIB [12] 19.7 21.0 21.3 19.7 20.5 25.0 34.4
Human-Thinking + IMIS-Net [6] 34.0 27.1 30.1 33.4 37.5 37.7 32.8
Human-Thinking + MediSee [39] 43.6 37.0 34.4 31.9 29.7 37.0 25.1
GPT-4o [1] + IMIS-Net [6] 25.1 20.4 18.1 17.0 11.6 12.5 18.9
Gemini-2.5-Pro [38] + IMIS-Net [6] 27.2 28.6 29.4 26.2 22.6 20.1 30.5
Qwen3-VL [3] + IMIS-Net [6] 25.6 24.1 20.3 21.8 15.2 14.8 27.6
GPT-4o [1] + MediSee [39] 37.7 33.4 27.3 25.5 14.9 21.5 20.9
Gemini-2.5-Pro [38] + MediSee [39] 30.8 30.4 29.5 26.7 31.5 21.5 33.3
Qwen3-VL [3] + MediSee [39] 37.7 33.9 30.1 25.4 20.2 34.7 29.9
SegLLM-7B [42] 35.2 39.5 27.2 30.1 34.7 37.4 34.1
SegLLM-13B [42] 36.4 42.1 32.5 33.2 35.3 39.6 36.0
MediRound + READ [30] 49.1 58.3 55.0 56.2 62.4 63.3 44.5
MediRound 50.2 60.8 55.9 58.8 63.7 64.7 46.1
MediRound + JCM 52.6 63.6 59.7 64.6 69.5 66.3 54.8
Table 3: Vanilla medical referring segmentation results.
Methods Metrics
Dice gIoU cIoU
Medical Interaction Seg. (bbox input)
MedSAM [25] 77.7 67.1 81.0
IMIS-Net [6] 77.0 67.5 87.2
Medical Referring Seg.
Grounding DINO [22] + SAM-Med2D [5] 18.4 12.5 10.9
Grounding DINO [22] + MedSAM [25] 18.5 11.0 13.5
BiomedParse [50] 24.9 20.1 14.5
MedPLIB [12] 40.6 39.1 45.2
IMIS-Net [6] 41.3 35.5 65.8
MediSee [39] 61.2 50.4 70.8
MediRound + READ [30] 60.7 48.0 65.2
MediRound 62.1 48.3 66.3
5 Experiments
5.1 Experimental Setting
Network Architecture.   For MediRound, as described in Sec. 4, we employ LLaVA-Med-v1.5-Mistral-7B [19] as the MLLM, and MedSAM [25] as the vision backbone. The $\mathcal{G}_{b}^{enc}$ is implemented as a linear projection from 4 dimensions to 4096. For JCM, the projections in the Quality Judgment Module and Correction Module are implemented as three-layer MLPs with dimensions [256, 512, 512, 1] and [256, 512, 512, 256], respectively.
Implementation Details.   MediRound is trained on the MR-MedSeg dataset using four NVIDIA RTX A6000 GPUs (48 GB memory each) for approximately 1.5 days. The batch size per device is set to 15. To enhance training efficiency, we adopt the DeepSpeed [33] engine. Optimization is performed using the AdamW [23] optimizer with a learning rate of 0.0003. Learning rate scheduling is managed by the WarmupDecayLR scheduler, with a warm-up phase spanning 100 iterations. The LoRA rank is uniformly set to 8. For training JCM, we keep the same hyperparameter settings, and it can be completed in just half a day. All training is conducted on the MR-MedSeg training set, employing a teacher forcing strategy [44].
5.2 Multi-Round Experimental Results
The results of multi-round medical reasoning segmentation are summarized in Tables 1 and 2. Table 1 reports the overall performance across all rounds, while Table 2 provides detailed scores for each individual round of interaction. Since existing text-prompt-based medical segmentation methods lack the capability for multi-round entity-level reasoning segmentation, we first incorporate human guidance to enable a broader and more equitable comparison. Specifically, humans with medical expertise assist traditional segmentation models through multi-round reasoning, allowing these models to perform the task. As shown in Tables 1 and 2, human assistance can substantially improve the model’s ability to perform multi-round segmentation. However, the overall performance remains suboptimal. This is primarily because, even with a certain level of medical knowledge, humans may still make errors when dealing with complex multi-round medical reasoning tasks, and their coordination with the model tends to be relatively rigid. In contrast, the proposed MediRound can effectively comprehend the current query while simultaneously integrating historical information and mask results from referred rounds. By achieving this capability, it demonstrates an average improvement of approximately 15% across various evaluation metrics compared to other methods. Notably, despite its lightweight design, the proposed JCM remarkably improves the model’s overall performance, particularly as the number of interaction rounds increases (refer to Table 2). This improvement is primarily attributed to the fact that JCM refines suboptimal mask outputs from earlier rounds before they are referred in subsequent ones, thereby effectively mitigating the error accumulation inherent in the chain-like pipeline of multi-round segmentation.
We further construct a hybrid paradigm for comparison, which integrates an MLLM with a medical segmentation model. Specifically, we leverage the strong text–image reasoning capabilities of MLLMs [3, 1, 38] to interpret complex multi-round queries, and subsequently feed their generated reasoning outputs into a medical segmentation model to produce visual predictions. As shown in Tables 1 and 2, despite being supported by powerful MLLMs, these combined approaches still underperform compared with MediRound. We attribute this advantage to our model’s end-to-end training on large-scale medical image–text pairs, which enables more coherent dynamic alignment between multimodal features than the two-stage combined paradigms. In addition, we include SegLLM [42]—the exploratory work on multi-round reasoning segmentation in natural images—in our comparison. The results indicate that while SegLLM performs well in natural images, it struggles when generalized to medical scenarios.
Figure 6: Qualitative comparison with different methods on medical multi-round interactive segmentation.
5.3 Vanilla Medical Referring Segmentation
To demonstrate MediRound’s effectiveness in conventional single-round referring medical segmentation, we evaluate it against several prior related methods on the SA-Med2D-20M [48] benchmark. As summarized in Table 3, our model achieves strong performance, remaining highly competitive in traditional medical image segmentation. For context, we also report the segmentation results of IMIS-Net [6] and MedSAM [25] using conventional bounding box guidance.
Table 4: Ablation study on the choices of vision and MLLM backbones. Experiments are conducted on the MR-MedSeg val set.
Settings Dice gIoU cIoU
Seg. Head Backbone
SAM (ViT-H) 46.7 33.1 37.2
SAM-Med2D 53.9 45.3 55.5
MedSAM (LoRA) 54.6 44.9 54.4
MedSAM 55.8 46.5 55.8
MLLM Backbone
LLaVA-v1.5 (LoRA) 53.4 42.1 53.3
LLaVA-Med (Frozen) 48.0 39.2 49.6
LLaVA-Med (LoRA) 55.8 46.5 55.8
Figure 7: Ablation study on the threshold $\beta$ value in JCM.
5.4 Qualitative Analysis
We present qualitative comparisons in Figure 6. Existing methods largely fail in multi-round reasoning for medical segmentation, whereas MediRound can understand the current query and dialogue history for accurate segmentation.
5.5 Ablation Study
Choices of $\bm{\beta}$ in JCM.   We investigate the impact of the threshold $\beta$ in the Quality Judgement Module of JCM. As shown in Figure 7, the collaboration between JCM and MediRound achieves the best performance when $\bm{\beta=0.6}$. This is likely because a smaller $\beta$ causes many suboptimal results to lose the opportunity for refinement, while an excessively large $\beta$ may instead disturb some high-quality predictions. For the definition and role of $\beta$, see Algorithm 1.
Choices of the Main Components.   Table 4 summarizes the effects of different foundational MLLMs and segmentation head backbones. The results show that LLaVA-Med achieves superior performance owing to its strong multimodal alignment capability in medical contexts, while the MedSAM performs best as the segmentation head, benefiting from large-scale medical image pretraining.
6 Conclusion
This work introduces MEMR-Seg, a new task that advances medical image segmentation toward interactive, multi-round reasoning. By constructing the large-scale MR-MedSeg dataset and developing the MediRound baseline with a helpful Judgment & Correction Mechanism, this work demonstrates the feasibility and effectiveness of medical multi-round entity-level reasoning, providing insights for future research on interactive medical segmentation.
7 Appendix
Abstract   Due to the space limit, we provide some implementation details and extra discussion in this Appendix. Specifically, the content of this document can be summarized as follows:
  • We present the training details of Judgment & Correction Mechanism (JCM) in Appendix A.
  • Additional training details of MediRound are provided in Appendix B.
  • We conduct ablation study on the effects of different strategies for extracting reference-mask information, as detailed in Appendix C.
  • We conduct further analyses of the proposed MR-MedSeg dataset in Appendix D.
  • Additional visualization examples are provided in Appendix E to illustrate MediRound and MR-MedSeg. You can also watch the MediRound demo video (MediRound_Demo.mp4) for a more intuitive understanding.
  • The limitations and future works of our paper are discussed in Appendix F.
Figure A: Overview of the JCM training pipeline. Please zoom in for improved clarity.
Appendix A Training the JCM
This section presents the training details of the Judgment & Correction Mechanism (JCM). The overall training pipeline of JCM is illustrated in Figure A.
A.1 Training Pipeline
The overall training pipeline of JCM is illustrated in Figure A. As shown in Figure A, LLaVA-Med takes $x_{all}$ (denoting the full set of inputs) and produces the [SEG] token as output. The hidden feature of [SEG], $h_{c}$, is first fed into the Correction Module $MLP_{C}$, producing $h^{\prime}_{c}$. The mask decoder $\mathcal{G}_{i}^{dec}$ then takes $h^{\prime}_{c}$ along with the visual features $f$ (extracted by the vision backbone $\mathcal{G}_{i}^{enc}$) and outputs the predicted mask $\hat{\mathbf{M}}$. In parallel, $\mathcal{G}_{i}^{dec}$ takes $h_{c}$, $f$ and outputs the uncorrected mask $\tilde{\mathbf{M}}$. The process can be formulated as:
$h^{\prime}_{c}={MLP}_{C}(h_{c}),\quad\hat{\mathbf{M}}=\mathcal{G}_{i}^{dec}(f,h^{\prime}_{c}),\quad\tilde{\mathbf{M}}=\mathcal{G}_{i}^{dec}(f,h_{c}).$ (1)
Additionally, we also feed $h_{c}$ into the Quality Judgment Module $MLP_{J}$, which outputs a quality score $q\in[0,1]$ for $h_{c}$:
$q=MLP_{J}(h_{c}),\quad q\in[0,1]$ (2)
We utilize the end-to-end trained MediRound weights to optimize the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ within the JCM. During the training of JCM, all parameters of the MediRound modules (e.g., LLaVA-Med, MedSAM, etc.) are frozen, and only the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ are trainable.
A.2 Optimization Strategy
For optimization, as illustrated in Figure A, the Correction Module $MLP_{C}$ and the Quality Judgment Module $MLP_{J}$ are jointly optimized using the mask loss $\mathcal{L}_{mask}$ and the quality score loss $\mathcal{L}_{score}$. For the quality score loss $\mathcal{L}_{score}$, we compute the Binary Cross-Entropy loss $\mathcal{L}_{bce}^{QJ}$ between the predicted quality score $q\in[0,1]$ and the ground-truth quality score $\mathbf{Q}\in[0,1]$. The ground-truth quality score $\mathbf{Q}$ is obtained by computing the Intersection-over-Union (IoU) between the uncorrected mask $\tilde{\mathbf{M}}$ and the ground-truth mask $\mathbf{M}$, serving as a soft target in the Binary Cross-Entropy loss $\mathcal{L}_{bce}^{QJ}$. The mask loss $\mathcal{L}_{mask}$ is also calculated between the predicted mask $\hat{\mathbf{M}}$ and the ground-truth mask $\mathbf{M}$, and it follows the same formulation as that used during MediRound training. The overall loss $\mathcal{L}_{JCM}$ can be formulated as:
[…]
Appendix B MediRound Training Details
B.1 Training Parameter
In end-to-end training for MediRound, to preserve the fundamental capabilities of the LLM in $\mathcal{G}_{i}$ and save computing resources, we leverage LoRA to perform efficient fine-tuning on the LLM in $\mathcal{G}_{i}$, while completely freezing the vision backbone $\mathcal{G}_{i}^{enc}$ and LLaVA-Med vision encoder $\mathcal{G}_{v}^{enc}$. The mask decoder $\mathcal{G}_{i}^{dec}$ and the bounding box encoder $\mathcal{G}_{b}^{enc}$ are fully trainable.
B.2 Optimization Strategy
MediRound is trained through two supervisions. As for text generation, we compute the Auto-Regressive Cross-Entropy loss $\mathcal{L}_{txt}$ between the generated text output $\hat{y}_{txt}$ and the ground-truth text answer $y_{txt}$. For the generation of high-quality segmentation masks, the mask loss $\mathcal{L}_{mask}$ is calculated between the predicted mask $\hat{\mathbf{M}}$ and the ground-truth mask $\mathbf{M}$. And the mask loss $\mathcal{L}_{mask}$ is a weighted sum of the per-pixel Binary Cross-Entropy $\mathcal{L}_{bce}$ and the DICE loss $\mathcal{L}_{dice}$, with the weights determined by $\lambda_{bce}$ and $\lambda_{dice}$. The overall loss $\mathcal{L}_{MediRound}$ is formulated as:
[…]
Table A: Ablation study on the effects of different strategies for extracting reference-mask information. The experimental results are reported on the evaluation set of MR-MedSeg. $\mathcal{C}$: Cropped image input; $\mathcal{B}$: Bounding box input
$\mathcal{C}$ $\mathcal{B}$ Dice gIoU cIoU
41.9 32.5 38.4
55.4 45.8 56.0
55.8 46.5 55.8
Appendix C Additional Ablation Study
We further investigate the impact of different strategies for extracting information from the reference mask. As shown in Table A, the experimental results indicate that using both the cropped image and the bounding box as reference-mask information yields the best performance.
Appendix D Additional Analyses of MR-MedSeg
We provide additional analyses and clarifications of the proposed MR-MedSeg dataset to more clearly characterize its properties and data distribution.
D.1 How to Refer to Previous Results
In MR-MedSeg dataset, we primarily employ three phrasing strategies to refer to previous round results. The first is “based on / … / referring to the output of round NN; the second is “based on / … / referring to instance NN. The third is more specific: when referring directly to the immediately preceding round and when grammatical structure allows, we may simply use “its …” to refer to the previous result.
D.2 Detailed Descriptions of Each Multi-Round Medical Segmentation Scenario
As illustrated in the Figure 2 (main text), we design five distinct types of interaction scenarios tailored for multi-round medical reasoning segmentation, each reflecting realistic use cases of interactive medical image segmentation:
  • Organ-Lesion: In this scenario, the inter-instance relationships within the data are characterized by the associations between organs (or tissues) and the lesions located in those regions. These data first prompt the model to segment individual organs or tissues, and subsequently require it to segment the lesions that occur on these previously segmented structures. Consequently, accurate lesion segmentation requires the model to establish a cross-round understanding of the correspondence between lesions and the organs segmented in prior rounds.
  • Anatomical Structure Stratification: In this scenario, the inter-instance relationships within the data correspond to the hierarchical anatomical relationships among different tissue structures. The model is first prompted to segment major anatomical regions, followed by the segmentation of substructures within these regions. Consequently, accurate segmentation of finer tissue structures necessitates cross-round understanding of the larger anatomical structures segmented in earlier rounds.
  • Organ/Tissue Attribute Relationship: In this scenario, the inter-instance relationships within the data correspond to the biological attribute associations among different organs or tissues. The data first prompts the model to segment a specific organ or tissue, and then, through the described attribute relationships, instructs it to segment the corresponding related structures. Consequently, accurate segmentation of medical entities in later rounds requires the model to reason over attribute relationships while maintaining cross-round understanding of previously segmented entities.
  • Spatial Relationship: The inter-instance relationships in this scenario within the data correspond to the relative spatial relationships among different organs or tissues in a medical image. The data first prompts the model to segment a particular structure and subsequently, through the relative positional descriptions based on the previously segmented structures, requires it to segment other related structures. Consequently, accurate segmentation of medical entities in later rounds necessitates both the ability to reason over spatial relationships and cross-round understanding of previously segmented entities.
  • Strong Inferential Relationship: Some medical images contain multiple medical entities of the same class. We transform such data to design a scenario characterized by Strong Inferential Relationship. Specifically, in the first round, we prompt the model to segment the initial entity through detailed referring expressions (e.g., segment the left lung or segment the first left rib). In subsequent rounds, we employ queries based on these strong inferential relationships to instruct the model to segment other instances of the same class (e.g., segment the other lung or segment the next left rib). For data of this scenario type, accurate segmentation of targets in subsequent rounds necessitates that the model maintain cross-round understanding of the entities segmented in earlier rounds, thereby requiring robust reasoning abilities across different segmentation rounds.
Table B: Data statistics of images, conversations, and maximum dialogue rounds across different multi-round interaction scenarios.
Scenario Type Training Set Validation Set Test Set
# of Convs # of Images Max Rounds # of Convs # of Images Max Rounds # of Convs # of Images Max Rounds
Organ-Lesion 39711 27908 5 428 374 5 431 371 5
Anat. Struct. Strat. 45910 42913 12 235 235 12 233 233 12
Spatial Relationship 40969 13313 2 196 196 2 196 196 2
Strong Inferential Rel. 38139 22369 12 260 226 12 262 227 12
Org/Tissue Attr. Rel. 10205 8883 7 151 98 7 151 101 7
Figure B: Analysis of the data structure in MR-MedSeg dataset. (a) Word cloud of MR-MedSeg; (b) Frequency distribution of each medical object across all the dialogues. Please zoom in for improved clarity.
D.3 Category Distribution
As described in the main text, MR-MedSeg comprises 168 distinct categories of medical entities. The word cloud presented in Figure B (a) visually represents the distribution of classes within our dataset, highlighting the diverse range of medical entity categories upon which our dialogues are based. We also conduct a comprehensive analysis of the frequency distribution of each medical entity category across all dialogues, as illustrated in Figure B (b). The results indicate that the category occurrence frequencies in MR-MedSeg follow an approximately long-tailed distribution, which aligns well with real-world medical scenarios.
Figure C: Additional interactive cases of MediRound.
D.4 Image Modality
As described in the main text, the proposed MR-MedSeg dataset encompasses a diverse set of imaging modalities. Specifically, it contains a total of 9 image modalities: CT, Endoscopy, Dermoscopy, X-ray, MR-FLAIR, MR-T1C/MR-T1CE, MR-T1, MR-T2, and MR-ADC. The diverse range of imaging modalities better simulates the complex variety of imaging types encountered in the real-world medical domain.
D.5 Data Statistics of Different Multi-Round Medical Segmentation Scenarios
To better characterize the data distribution across different scenarios, in Table B, we report the number of images and conversations for each multi-round interaction scenario, as well as the maximum dialogue rounds contained within each scenario.
Appendix E More Visualization Examples
E.1 Additional Interactive Cases
We provide additional interactive cases of MediRound in Figure C. A demonstration video of our model, designated “MediRound_Demo.mp4”, has also been included in the supplementary materials to provide a more detailed understanding.
Appendix F Limitations and Future Works
In this section, we examine the limitations of the present work and outline potential directions for future research.
F.1 Multi-Object Segmentation in One Round
The current MediRound configuration restricts each round to segmenting a single medical entity, thereby preventing the model from handling multiple objects within the same interaction turn. This limitation may introduce inconvenience in certain practical clinical scenarios. Accordingly, in future work, we aim to extend MediRound to support multi-target segmentation within each dialogue round.
F.2 Processing Long Sequence
We observe that excessively long queries in each round can negatively affect the segmentation accuracy of MediRound. This reduced robustness to long sequences may be attributed to the inherent sequence-length limitations of transformer-based architectures.
F.3 Overfitting in the Segmentation Tasks
Trained on large-scale segmentation data, MediRound becomes specialized for segmentation tasks and nearly completely loses the core conversational capabilities of the MLLM. This reflects a common phenomenon of task-specific overfitting observed when MLLMs are adapted to downstream tasks. In future works, we aim to address this challenge by enabling the model to maintain a certain level of medical vision–language dialogue ability while still performing segmentation accurately.
F.4 More Modalities
Currently, MediRound is unable to process three-dimensional medical data or medical videos, which are frequently encountered in real-world clinical scenarios. In future work, we plan to extend MediRound to additional modalities, thereby overcoming its current restriction to two-dimensional images.